# Unraveling the Key Components of OOD Generalization via Diversification

### v2

@@ Abstract
Supervised learning datasets may contain multiple cues that explain the training set equally well, i.e., learning any of them would lead to the correct predictions on the training data. However, many of them can be spurious, i.e., lose their predictive power under a distribution shift and consequently fail to generalize to out-of-distribution (OOD) data. Recently developed "diversification" methods (Lee et al., 2023; Pagliardini et al., 2023) approach this problem by finding multiple diverse hypotheses that rely on different features. This paper aims to study this class of methods and identify the key components contributing to their OOD generalization abilities. We show that (1) diversification methods are highly sensitive to the distribution of the unlabeled data used for diversification and can underperform significantly when away from a method-specific sweet spot. (2) Diversification alone is insufficient for OOD generalization. The choice of the used learning algorithm, e.g., the model’s architecture and pretraining, is crucial. In standard experiments (classification on Waterbirds and Office-Home datasets), using the second-best choice leads to an up to 20% absolute drop in accuracy. (3) The optimal choice of learning algorithm depends on the unlabeled data and vice versa i.e. they are co-dependent. (4) Finally, we show that, in practice, the above pitfalls cannot be alleviated by increasing the number of diverse hypotheses, the major feature of diversification methods. These findings provide a clearer understanding of the critical design factors influencing the OOD generalization abilities of diversification methods. They can guide practitioners in how to use the existing methods best and guide researchers in developing new, better ones. footnotetext: $*{}^{*}$qual contribution. Corresponding author: harold.benoit@alumni.epfl.ch
@@ 1 Introduction
Figure 1: Diversification is a two-legged problem where unlabeled data and learning algorithm both matter and are co-dependent. Refer to caption and Refer to caption represent the training data points and their labels. Refer to caption represents unlabeled data. $h_{\mathrm{ERM}}$ represents the hypothesis found by empirical risk minimization (ERM), thus, reflecting the inductive bias of the learning algoritm. $h_{2}$ represents a second diverse hypothesis found by a diversification method; it has low error training data as $h_{\mathrm{ERM}}$ does, but disagrees with it on the unlabeled data. Compared to the original setting in (a), in this work, we study how changing (b) unlabeled data, and (c) the learning algorithm lead to different solutions and, therefore, performance.
@@ 2 Related Work
Spurious correlation and underspecification. As a special case of OOD generalization problem, spurious correlations can arise from the underspecified nature of the training data (D’Amour et al., 2020; Koh et al., 2021; Liang & Zou, 2022). In this setting, neural networks tend to learn simple (spurious) concepts rather than the true causal concepts, a phenomenon known as simplicity bias (Shah et al., 2020; Huh et al., 2023) or shortcut learning (Geirhos et al., 2020; Scimeca et al., 2022). Some works combat spurious correlation by improving worst-group performance (Sagawa et al., 2020; Hu et al., 2018; Zhang et al., 2021), some require group annotation (Sagawa et al., 2020; Zhang et al., 2022; Creager et al., 2021) and others (Liu et al., 2021; LaBonte et al., 2022; Sohoni et al., 2020) aim at the no group information scenario. Diversification methods fit into the latter, as they only rely on additional unlabeled data to promote diversity between multiple hypotheses.
Diversification methods. Recently proposed diversification methods (Lee et al., 2023; Pagliardini et al., 2023; Teney et al., 2022a, b) find multiple diverse hypotheses during training to handle spurious correlations. They introduce an additional diversification loss over multiple trained hypotheses, forcing them to rely on different features while still fitting training data well. (Teney et al., 2022a, b) use input-space diversification that minimizes the alignment of input gradients over pairs of models at all training data points. DivDis (Lee et al., 2023) and D-BAT (Pagliardini et al., 2023) use output-space diversification, minimizing the agreement between models’ predictions on additional unlabeled data. We focus on studying the latter, as these methods outperform the input-space ones by a large margin, achieving state-of-the-art performance in the setting where true labels are close to or completely correlated with spurious attributes.
Inductive biases in learning algorithms. In this work, we study the influence of the choice of the learning algorithm and, hence, its inductive bias on the performance of diversification methods. Different learning algorithms have different inductive biases (Shalev-Shwartz & Ben-David, 2014; Hüllermeier et al., 2013), which makes a given algorithm to prioritize a specific solutions (Gunasekar et al., 2018; Ji & Telgarsky, 2020). While being highly overparameterized (Allen-Zhu et al., 2019) and able to fit even random labels (Zhang et al., 2017), deep learning models were shown to benefit from architectural (Xu et al., 2021; Naseer et al., 2021) , optimization (Kalimeris et al., 2019; Liu et al., 2020) and pre-training (Immer et al., 2022; Lovering et al., 2021) inductive biases. In our study we show that diversification methods are sensitive to the choice of architecture and pretraining method.
@@ 3.2 Diversification for OOD Generalization
DivDis (Lee et al., 2023) and D-BAT (Pagliardini et al., 2023) focus on the spurious correlation setting. They assume access to additional unlabeled data $D_{u}$ to find multiple diverse hypotheses that all fit the training data $(D_{t},h^{*})$ but disagree, i.e., make diverse predictions, on $D_{u}$. The motivation is to better cover the space $\mathcal{H}_{t}^{*}$ and, consequently, find a hypothesis from $\mathcal{H}^{*}_{t}\cap\mathcal{H}^{*}_{\mathrm{ood}}$ that also generalizes to OOD data.
Optimization objective. Following Lee et al. (2023); Pagliardini et al. (2023), we define a diversification loss $\mathnormal{A}_{D_{u}}(h_{1},h_{2})$ that quantifies the agreement between two hypotheses on $D_{u}$. Then, in the case of finding $K$ hypotheses, the training objective of a diversification method is the sum of ERM loss and the diversification loss averaged over all pairs of hypotheses:
@@
Diversification loss. Let $P_{h_{i}}$ be the predictive distribution of a hypothesis $h_{i}$ on given data $D$. We consider the following two diversification losses:
  • DivDis (Lee et al., 2023): $\mathnormal{A}_{D}(h_{1},h_{2})=D_{\mathrm{KL}}(P_{(h_{1},h_{2})}||P_{h_{1}}\otimes P_{h_{2}})+\lambda\sum_{i\in\{1,2\}}D_{\mathrm{KL}}(P_{h_{i}}||\hat{P}),$
    where the first term is the mutual information, which is equal to 0 iff $P_{h_{1}}$ and $P_{h_{2}}$ are independent. The second term is the KL-divergence between the predictive distribution of $h_{i}$ and a prior distribution $\hat{P}$, which is usually set to the distribution of labels in $D_{t}$. It prevents hypotheses from collapsing to degenerate solutions, such as predicting the same label for all samples.
@@
The two-stage framework. After finding $K$ hypotheses, one needs to be chosen to make the final prediction, leading to a two-stage approach (Lee et al., 2023), summarised as follows:
  1. 1.
    Diversification: find $K$ diverse hypotheses $\mathcal{H}_{K}^{*}\subset\mathcal{H}^{*}_{t}$.
  2. 2.
    Disambiguation: select one hypothesis $\hat{h}\in\mathcal{H}^{*}_{K}$ given additional information (e.g., a few test labeled examples or human supervision).
We identify the first diversification stage as the most critical one. Indeed, if the desired hypothesis $h^{*}$ is not chosen ($h^{*}\notin\mathcal{H}_{K}^{*}$), the second stage cannot make up for it as it is limited to only hypotheses from $\mathcal{H}_{K}^{*}$. We, therefore, focus on studying the first stage and assume access to the oracle that chooses the best available hypothesis in the second stage.
@@ 4.1 Theoretical and Empirical Study of a Synthetic Example
Figure 2: Performance of diversification is highly dependent on unlabeled OOD data. Left: Top-left quadrant: The 2D binary classification task. Other quadrants: Show the second hypotheses (arrows are normal vectors) found by D-BAT ($h_{2}^{\mathrm{DB}})$ and DivDis-Seq ($h_{2}^{\mathrm{DD}})$ with varied spurious ratios of unlabeled OOD data $r_{D_{u}}=\{0,0.25,0.5\}$ (from inversely correlated to balanced). Right: Best hypothesis test accuracy of D-BAT & DivDis(-Seq) on MNIST-CIFAR (M/C) for varied spurious ratios $r_{D_{u}}$ and number of hypotheses $K$. The test accuracy is measured on hold-out balanced data $D_{\mathrm{ood}}$ (i.e. $r_{D_{\mathrm{ood}}}=0.5$, no spurious correlation).
Synthetic 2D binary classification task. In Fig. 2-Left, we show a 2D task with distribution $D_{\mathrm{ood}}$ spanning a 2D square, i.e., $\{x=(x_{1},x_{2})\in[-1,1]^{2}\}$. We define our hypothesis space $\mathcal{H}$ to be all possible linear classifiers $h(x;\beta)$ where $\beta$ is the radian of the classification plane w.r.t horizontal axis $x_{1}$. The ground truth labeling function is defined as $h^{\star}(x)=h(x;\frac{\pi}{2})=\mathcal{I}\{x_{1}>0\}$ where $\mathcal{I}$ is the indicator function, and the training distribution is defined as $D_{t}=\{x=(x_{1},x_{2})\in\{[-1,0]\times[0,1]\}\cup\{[0,1]\times[-1,0]\}\}$. We then define a spurious feature function as $h_{\mathrm{sp}}(x)=h(x;0)=\mathcal{I}\{x_{2}<0\}$ and assume that ERM converges to $h_{\mathrm{sp}}$. This means that the first hypothesis $h_{1}^{DB}$ (D-BAT) and $h_{1}^{DD}$ (DivDis-Seq) of both methods converge to $h_{\mathrm{sp}}$.
Finally, we define different distributions of unlabeled data $D_{u}$ to have different spurious ratios $r_{D_{u}}$ from 0 to 0.5 (the construction is described in Appendix A).
@@ 5.1 Diversification Alone Is Insufficient for OOD Generalization
Notably, we prove that having as many diverse hypotheses as the number of data points in $D_{\mathrm{ood}}$ is still insufficient to guarantee better than a random guess accuracy. Indeed, there always exists a set of hypotheses satisfying all the constraints of the diversification objective in Eq. 2 while having random accuracy w.r.t the true labeling $h^{*}$ on OOD data. The following Proposition 2 formalizes this intuition in the binary classification case. Please see its proof and extension to multi-class case in Appendix E.
@@ 5.2 Learning Algorithm Selection: A Key to Effective Diversification
Figure 3: The performance of diversification methods is highly sensitive to the choice of architecture and pretraining method. Left: DivDis and D-BAT best hypothesis ($K=2$) performance with multiple pretraining strategies and architecture pairs on Waterbirds-CC (Left) and Office-Home (Right). ResNet50 is used, if not specified. Right: Top-1 accuracy on ImageNet-1k after fine-tuning.
@@
Waterbirds-CC D-BAT (ViT-B/16 IN) 57.1±${}_{\pm 3.7}$ 57.1±${}_{\pm 3.7}$ 57.1±${}_{\pm 3.7}$ 57.1±${}_{\pm 3.7}$
DivDis (MoCo-v2) 49.4±${}_{\pm 10.3}$ 51.7±${}_{\pm 6.0}$ 49.6±${}_{\pm 8.3}$ 48.4±${}_{\pm 0.9}$
Office-Home D-BAT (ViT-MAE) 61.9±${}_{\pm 0.7}$ 62.6±${}_{\pm 0.1}$ 62.6±${}_{\pm 0.1}$ 62.6±${}_{\pm 0.1}$
DivDis (Resnet50 IN) 55.9±${}_{\pm 0.6}$ 54.6±${}_{\pm 0.1}$ 53.6±${}_{\pm 0.4}$ 53.1±${}_{\pm 0.2}$
Table 1: Increasing the number of hypotheses does not bridge the performance gap between different models. We increase the number of hypotheses found by diversification methods for the second-best model in Fig. 3 and find that it is not enough to bridge the performance gap with the best-performing model. Best hypothesis accuracy is reported. Results are averaged over 3 seeds. Standard deviations for Waterbirds-CC are larger due to the usage of the worst-group accuracy metric.
Increasing $K$ does not improve performance. Finally, we study whether the performance gap between the best and second-best models tested in Fig. 3 can be closed by increasing the number of hypotheses $K$, as this is allegedly the major feature and motivation of diversification methods. Tab. 1 shows that, similar to the observation made in Sec. 4.2, increasing $K$ does not bring any improvements, suggesting that the choice of the model is more important for enabling OOD generalization. In Fig. 14, we further show that DivDis does not scale well to larger $K$ (e.g., $K=64$) “out-of-the-box”, and the performance drops as the number of hypotheses increases. Note that testing D-BAT in this regime would be prohibitively expensive.
Takeaway. Diversification methods are highly sensitive to the choice of the learning algorithm, e.g., architecture and pretraining method. The “built-in” mechanism of increasing the number of hypotheses $K$ does not alleviate this issue and fails at improving performance.
@@ 5.3 On the Co-Dependence between Learning Algorithm and Unlabeled Data
Sec. 4 and Sec. 5.2 show the sensitivity of the diversification methods to the distribution of the unlabeled data and the choice of a learning algorithm respectively. Here we further demonstrate that these choices are co-dependent, i.e., the optimal choice for one depends on the other. Specifically, we show that by only varying the distribution of unlabeled data, the optimal architecture can be changed.
@@
As the true labeling $h^{*}$, we use a binary classification task constructed by splitting the original 10 classes into two sets of five. Then, as Tab. 2-Right illustrates, we construct training data $D_{t}$ to contain samples where all $h^{*}$, $h_{\mathrm{MLP}}$, and $h_{\mathrm{RN}}$ agree, i.e., $D_{t}=\{x\in D:h^{*}(x)=h_{\mathrm{MLP}}(x)=h_{\mathrm{RN}}(x)\}$. Thus, by design, both $h_{\mathrm{MLP}}$ and $h_{\mathrm{RN}}$ are completely spuriously correlated with $h^{*}$. Then, we define unlabeled OOD data $D_{u}$ s.t. either $(r^{h_{\mathrm{MLP}}}_{D_{u}}=0,r^{h_{\mathrm{RN}}}_{D_{u}}=\frac{1}{2})$ (denoted as $h^{*}\perp h_{\mathrm{RN}}$), or $(r^{h_{\mathrm{MLP}}}_{D_{u}}=\frac{1}{2},r^{h_{\mathrm{RN}}}_{D_{u}}=0)$ (denoted as $h^{*}\perp h_{\mathrm{MLP}}$).
This means that $h^{*}$ is inversely correlated with only one of $h_{\mathrm{MLP}}$ or $h_{\mathrm{RN}}$, while not correlated to the other hypothesis.
Results. Keeping the training data fixed, we train D-BAT ($K=2$) using different architecture and constructed unlabeled data pairs $(\mathcal{A},D_{u})$. Tab. 2-Left shows that the performance of $\mathcal{A}$ drops to almost random chance when $h_{\mathcal{A}}$ does not inversely correlate with $h^{*}$ on the unlabelled data (② and ③). This is consistent with Sec. 4, where we show that the setting with $r^{h_{1}}_{D_{u}}=\frac{1}{2}$ is disadvantageous for D-BAT. In Appendix H, Tab. 11 further shows similar observation for a different architecture pair (ViT & ResNet18), and Fig. 15 extends the experiment with a smooth interpolation from one unlabeled dataset setting to the other, showing a linear transition where one architecture goes from optimal performance to random-chance accuracy, and vice-versa.
@@
$D_{u}$ $\mathcal{A}$ Test Acc.(%) $h^{\star}\perp h_{\mathrm{MLP}}$ MLP 89.2±${}_{\pm 0.8}$ ResNet18 56.7±${}_{\pm 0.8}$ $h^{\star}\perp h_{\mathrm{RN}}$ MLP 55.4±${}_{\pm 0.3}$ ResNet18 77.0±${}_{\pm 0.7}$ [Uncaptioned image]
Table 2: The optimal choices of the (learning algorithm) architecture and unlabeled data are co-dependent. Left: Performance of the D-BAT method using two architectures and two created unlabeled data distributions. Test accuracy is on hold-out data $D_{\mathrm{ood}}\sim D_{u}$. Right: Illustration of the two unlabeled data distributions used in the experiment. We keep the training data fixed and change $D_{u}$ s.t. $h^{\star}$ is inversely correlated (denoted by $\perp$’) to one of the “spurious” hypotheses $h_{\mathrm{MLP}},h_{\mathrm{RN}}$. We show that the optimal choice of the architecture ($\mathcal{A}$) depends on the unlabeled data distribution, and the best $\mathcal{A}$ is the one for which $h^{*}\perp h_{\mathcal{A}}$ on $D_{u}$.

### v3

@@ Abstract
Supervised learning datasets may contain multiple cues that explain the training set equally well, i.e., learning any of them would lead to the correct predictions on the training data. However, many of them can be spurious, i.e., lose their predictive power under a distribution shift and consequently fail to generalize to out-of-distribution (OOD) data. Recently developed "diversification" methods (Lee et al., 2023; Pagliardini et al., 2023) approach this problem by finding multiple diverse hypotheses that rely on different features. This paper aims to study this class of methods and identify the key components contributing to their OOD generalization abilities. We show that (1) diversification methods are highly sensitive to the distribution of the unlabeled data used for diversification and can underperform significantly when away from a method-specific sweet spot. (2) Diversification alone is insufficient for OOD generalization. The choice of the used learning algorithm, e.g., the model’s architecture and pretraining, is crucial. In standard experiments (classification on Waterbirds and Office-Home datasets), using the second-best choice leads to an up to 20% absolute drop in accuracy. (3) The optimal choice of learning algorithm depends on the unlabeled data and vice versa, i.e., they are co-dependent. (4) Finally, we show that, in practice, the above pitfalls cannot be alleviated by increasing the number of diverse hypotheses, the major feature of diversification methods. These findings provide a clearer understanding of the critical design factors influencing the OOD generalization abilities of diversification methods. They can guide practitioners in how to use the existing methods best and guide researchers in developing new, better ones. footnotetext: Equal contribution. Corresponding author: harold.benoit@alumni.epfl.ch
@@ 1 Introduction
Figure 1: Diversification is a two-legged problem where unlabeled data and learning algorithm both matter and are co-dependent. Refer to caption and Refer to caption represent the training data points and their labels. Refer to caption represents unlabeled data. $h_{\mathrm{ERM}}$ represents the hypothesis found by empirical risk minimization (ERM), thus reflecting the inductive bias of the learning algorithm. $h_{2}$ represents a second diverse hypothesis found by a diversification method; it has low error on training data as $h_{\mathrm{ERM}}$ does, but disagrees with it on the unlabeled data. Compared to (a) the original setting, we study how changing (b) unlabeled data and (c) the learning algorithm yield different solutions and, therefore, performance.
@@ 2 Related Work
Spurious correlation and underspecification. As a special case of OOD generalization problem, spurious correlations can arise from the underspecified nature of the training data (D’Amour et al., 2020; Koh et al., 2021; Liang & Zou, 2022). In this setting, neural networks tend to learn simple (spurious) concepts rather than the true causal concepts, a phenomenon known as simplicity bias (Shah et al., 2020; Huh et al., 2023) or shortcut learning (Geirhos et al., 2020; Scimeca et al., 2022). Some works combat spurious correlation by improving worst-group performance (Sagawa et al., 2020; Hu et al., 2018; Zhang et al., 2021), some require group annotation (Sagawa et al., 2020; Kirichenko et al., 2022; Goel et al., 2021) and others (Liu et al., 2021; LaBonte et al., 2022; Sohoni et al., 2020; Zhang et al., 2022; Creager et al., 2021) aim at the no group information scenario. Diversification methods fit into the latter, as they only rely on additional unlabeled data.
Diversification methods. Recently proposed diversification methods (Lee et al., 2023; Pagliardini et al., 2023; Teney et al., 2022a; b) find multiple diverse hypotheses during training to handle spurious correlations. They introduce an additional diversification loss over multiple trained hypotheses, forcing them to rely on different features while still fitting training data well. Teney et al. (2022a; b) use input-space diversification that minimizes the alignment of input gradients over pairs of models at all training data points. DivDis (Lee et al., 2023) and D-BAT (Pagliardini et al., 2023) use output-space diversification, minimizing the agreement between models’ predictions on additional unlabeled data. We focus on studying the latter, as these methods outperform the input-space ones by a large margin, achieving state-of-the-art performance in the setting where true labels are close to or completely correlated with spurious attributes.
Inductive biases in learning algorithms. Different learning algorithms have different inductive biases (Shalev-Shwartz & Ben-David, 2014; Hüllermeier et al., 2013), which makes a given algorithm prioritize a specific solution (Gunasekar et al., 2018; Ji & Telgarsky, 2020). While being highly overparameterized (Allen-Zhu et al., 2019) and able to fit even random labels (Zhang et al., 2017), deep learning models were shown to benefit from architectural (Xu et al., 2021; Naseer et al., 2021) , optimization (Kalimeris et al., 2019; Liu et al., 2020) and pre-training (Immer et al., 2022; Lovering et al., 2021) inductive biases. In this work, we study the influence of the choice of the learning algorithm and, hence, its inductive bias on the performance of diversification methods. We show that diversification methods are sensitive to the choice of architecture and pretraining method.
@@ 3.2 Diversification for OOD Generalization
DivDis and D-BAT focus on the above spurious correlation setting. They assume access to additional unlabeled data $D_{u}$ to find multiple diverse hypotheses that all fit the training data $(D_{t},h^{*})$ but disagree, i.e., make diverse predictions, on $D_{u}$. The motivation is to better cover the space $\mathcal{H}_{t}^{*}$ and, consequently, find a hypothesis from $\mathcal{H}^{*}_{t}\cap\mathcal{H}^{*}_{\mathrm{ood}}$ that also generalizes to OOD data.
Optimization objective. Following DivDis and D-BAT, we define a diversification loss $\mathnormal{A}_{D_{u}}(h_{1},h_{2})$ that quantifies the agreement between two hypotheses on $D_{u}$. Then, in the case of finding $K$ hypotheses, the training objective of a diversification method is the sum of ERM loss and the diversification loss averaged over all pairs of hypotheses:
@@
Diversification loss. Let $P_{h_{i}}$ be the predictive distribution of a hypothesis $h_{i}$ on given data $D$. Then,
  • DivDis (Lee et al., 2023): $\mathnormal{A}_{D}(h_{1},h_{2})=D_{\mathrm{KL}}(P_{(h_{1},h_{2})}||P_{h_{1}}\otimes P_{h_{2}})+\lambda\sum_{i\in\{1,2\}}D_{\mathrm{KL}}(P_{h_{i}}||\hat{P}).$ The first term is the mutual information, which is equal to 0 iff $P_{h_{1}}$ and $P_{h_{2}}$ are independent. The second term is the KL-divergence between $P_{h_{i}}$ and a prior distribution $\hat{P}$ (usually set to the distribution of labels in $D_{t}$). It prevents hypotheses from collapsing to degenerate solutions.
@@
The two-stage framework. After finding $K$ hypotheses, one needs to be chosen to make the final prediction, leading to a two-stage approach (Lee et al., 2023): 1. Diversification: find $K$ diverse hypotheses $\mathcal{H}_{K}^{*}\subset\mathcal{H}^{*}_{t}$. 2. Disambiguation: select one hypothesis $\hat{h}\in\mathcal{H}^{*}_{K}$ given additional information (e.g., a few labeled test examples or human supervision). We identify the diversification stage as the most critical one. Indeed, if the desired hypothesis $h^{*}$ is not chosen ($h^{*}\notin\mathcal{H}_{K}^{*}$), the second stage cannot make up for it as it is limited to only hypotheses from $\mathcal{H}_{K}^{*}$. We, therefore, focus on the first stage and assume an oracle that chooses the best available hypothesis in the second stage.
@@ 4.1 Theoretical and Empirical Study of a Synthetic Example
Refer to captionRefer to caption
Figure 2: Performance of diversification is highly dependent on unlabeled OOD data. Left: Top-left quadrant: The 2D binary classification task. Other quadrants: Show the second hypotheses (arrows are normal vectors) found by D-BAT ($h_{2}^{\mathrm{DB}})$ and DivDis-Seq ($h_{2}^{\mathrm{DD}})$ with varied spurious ratios of unlabeled OOD data $r_{D_{u}}=\{0,0.25,0.5\}$ (from inversely correlated to balanced). Right: Best hypothesis test accuracy of D-BAT & DivDis(-Seq) on MNIST-CIFAR (M/C) for varied spurious ratios $r_{D_{u}}$ and number of hypotheses $K$. The test accuracy is measured on hold-out balanced data $D_{\mathrm{ood}}$ (i.e., $r_{D_{\mathrm{ood}}}=0.5$, no spurious correlation).
Synthetic 2D binary classification task. In Fig. 2-Left, we show a 2D task with distribution $D_{\mathrm{ood}}$ spanning a 2D square, i.e., $\{x=(x_{1},x_{2})\in[-1,1]^{2}\}$. We define our hypothesis space $\mathcal{H}$ to be all possible linear classifiers $h(x;\beta)$ where $\beta$ is the radian of the classification plane w.r.t horizontal axis $x_{1}$. The ground truth labeling function is defined as $h^{\star}(x)=h(x;\frac{\pi}{2})=\mathcal{I}\{x_{1}>0\}$ where $\mathcal{I}$ is the indicator function, and the training distribution is defined as $D_{t}=\{x=(x_{1},x_{2})\in\{[-1,0]\times[0,1]\}\cup\{[0,1]\times[-1,0]\}\}$. We then define a spurious feature function as $h_{\mathrm{sp}}(x)=h(x;0)=\mathcal{I}\{x_{2}<0\}$ and assume that ERM converges to $h_{\mathrm{sp}}$. This means that the first hypothesis $h_{1}^{DB}$ (D-BAT) and $h_{1}^{DD}$ (DivDis-Seq) of both methods converge to $h_{\mathrm{sp}}$. Finally, we define different distributions of unlabeled data $D_{u}$ to have different spurious ratios $r_{D_{u}}$ from 0 to 0.5 (the construction is described in Appendix A).
@@ 5.1 Diversification Alone Is Insufficient for OOD Generalization
Notably, we prove that having as many diverse hypotheses as the number of data points in $D_{\mathrm{ood}}$ is still insufficient to guarantee better than a random guess accuracy. Indeed, there always exists a set of hypotheses satisfying all the constraints of the diversification objective in Eq. 2 while having random accuracy w.r.t the true labeling $h^{*}$ on OOD data. The following Proposition 2 formalizes this intuition in the binary classification case. Please see its proof and extension to the multi-class case in Appendix E.
@@ 5.2 Learning Algorithm Selection: A Key to Effective Diversification
Refer to captionRefer to caption
Figure 3: The performance of diversification methods is highly sensitive to the choice of architecture and pretraining method. Left: DivDis and D-BAT best hypothesis ($K=2$) performance with multiple pretraining strategies and architecture pairs on Waterbirds-CC (Left) and Office-Home (Right). ResNet50 is used if not specified. Right: Top-1 accuracy on ImageNet-1k after fine-tuning.
@@
Waterbirds-CC D-BAT (ViT-B/16 IN) 57.1±3.7 57.1±3.7 57.1±3.7 57.1±3.7
DivDis (MoCo-v2) 49.4±10.3 51.7±6.0 49.6±8.3 48.4±0.9
Office-Home D-BAT (ViT-MAE) 61.9±0.7 62.6±0.1 62.6±0.1 62.6±0.1
DivDis (Resnet50 IN) 55.9±0.6 54.6±0.1 53.6±0.4 53.1±0.2
Table 1: Increasing the number of hypotheses does not bridge the performance gap between different models. We increase the number of hypotheses found by diversification methods for the second-best model in Fig. 3 and find that it is not enough to bridge the performance gap with the best-performing model. The best hypothesis accuracy is reported. Results are averaged over 3 seeds. Standard deviations for Waterbirds-CC are larger due to the usage of the worst-group accuracy metric.
Increasing $K$ does not improve performance. Finally, we study whether the performance gap between the best and second-best models tested in Fig. 3 can be closed by increasing the number of hypotheses $K$, as this is allegedly the major feature and motivation of diversification methods. Tab. 1 shows that similar to the observation made in Sec. 4.2, increasing $K$ does not bring any improvements, suggesting that the choice of the model is more important for enabling OOD generalization. In Fig. 14, we further show that DivDis does not scale well to larger $K$ (e.g., $K=64$) “out-of-the-box”, and the performance drops as the number of hypotheses increases. Note that testing D-BAT in this regime would be prohibitively expensive.
Takeaway. Diversification methods are highly sensitive to the choice of the learning algorithm, e.g., architecture and pretraining method. The “built-in” mechanism of increasing the number of hypotheses $K$ does not alleviate this issue and fails to improve performance.
@@ 5.3 On the Co-Dependence between Learning Algorithm and Unlabeled Data
Sec. 4 and Sec. 5.2 show the sensitivity of the diversification methods to the distribution of the unlabeled data and the choice of a learning algorithm, respectively. Here, we further demonstrate that these choices are co-dependent, i.e., the optimal choice for one depends on the other. Specifically, we show that by only varying the distribution of unlabeled data, the optimal architecture can be changed.
@@
As the true labeling $h^{*}$, we use a binary classification task constructed by splitting the original 10 classes into two sets of five. Then, as Tab. 2-Right illustrates, we construct training data $D_{t}$ to contain samples where all $h^{*}$, $h_{\mathrm{MLP}}$, and $h_{\mathrm{RN}}$ agree, i.e., $D_{t}=\{x\in D:h^{*}(x)=h_{\mathrm{MLP}}(x)=h_{\mathrm{RN}}(x)\}$. Thus, by design, both $h_{\mathrm{MLP}}$ and $h_{\mathrm{RN}}$ are completely spuriously correlated with $h^{*}$. Then, we define unlabeled OOD data $D_{u}$ s.t. either $(r^{h_{\mathrm{MLP}}}_{D_{u}}=0,r^{h_{\mathrm{RN}}}_{D_{u}}=\frac{1}{2})$ (denoted as $h^{*}\perp h_{\mathrm{RN}}$), or $(r^{h_{\mathrm{MLP}}}_{D_{u}}=\frac{1}{2},r^{h_{\mathrm{RN}}}_{D_{u}}=0)$ (denoted as $h^{*}\perp h_{\mathrm{MLP}}$). This means that $h^{*}$ is inversely correlated with only one of $h_{\mathrm{MLP}}$ or $h_{\mathrm{RN}}$, while not correlated to the other hypothesis.
Results. Keeping the training data fixed, we train D-BAT ($K=2$) using different architecture and construct unlabeled data pairs $(\mathcal{A},D_{u})$. Tab. 2-Left shows that the performance of $\mathcal{A}$ drops to almost random chance when $h_{\mathcal{A}}$ does not inversely correlate with $h^{*}$ on the unlabelled data (② and ③). This is consistent with Sec. 4, where we show that the setting with $r^{h_{1}}_{D_{u}}=\frac{1}{2}$ is disadvantageous for D-BAT. In Appendix H, Tab. 11 further shows similar observation for a different architecture pair (ViT & ResNet18), and Fig. 15 extends the experiment with smooth interpolation from one unlabeled dataset setting to the other, showing a linear transition where one architecture goes from optimal performance to random-chance accuracy, and vice-versa.
@@
$D_{u}$ $\mathcal{A}$ Test Acc.(%)
$h^{\star}\perp h_{\mathrm{MLP}}$ MLP 89.2±0.8
ResNet18 56.7±0.8
$h^{\star}\perp h_{\mathrm{RN}}$ MLP 55.4±0.3
ResNet18 77.0±0.7
Table 2: The optimal choices of the (learning algorithm) architecture and unlabeled data are co-dependent. Left: Performance of the D-BAT method using two architectures and two created unlabeled data distributions. The test accuracy is on hold-out data $D_{\mathrm{ood}}\sim D_{u}$. Right: Illustration of the two unlabeled data distributions used in the experiment. We keep the training data fixed and change $D_{u}$ s.t. $h^{\star}$ is inversely correlated (denoted by $\perp$’) to one of the “spurious” hypotheses $h_{\mathrm{MLP}},h_{\mathrm{RN}}$. We show that the optimal choice of the architecture ($\mathcal{A}$) depends on the unlabeled data distribution, and the best $\mathcal{A}$ is the one for which $h^{*}\perp h_{\mathcal{A}}$ on $D_{u}$.